# Almost All Vectorial Functions Have Trivial Extended-Affine Stabilizers

### 10 Mar 2026 v2

@@ 1 Introduction
Vectorial Boolean functions play a crucial role in symmetric cryptography, serving as S-boxes in block ciphers and hash functions. The classification of these functions is typically studied modulo certain equivalence relations that preserve important cryptographic properties. Extended-affine (EA) equivalence is one of the most fundamental such relations. It preserves differential uniformity, a key measure of resistance against differential cryptanalysis, as well as the Walsh spectrum and algebraic degree. A closely related notion is CCZ-equivalence, introduced by Carlet, Charpin, and Zinoviev [4], which also preserves some cryptographic properties and is more general than EA-equivalence.
A substantial body of recent work has focused on the algorithmic and structural aspects of EA- and CCZ-equivalence. Canteaut and Perrin [3] showed that CCZ-equivalence can be decomposed into EA-equivalence and an operation called twisting, and derived upper bounds on the number of EA-equivalence classes within a single CCZ-equivalence class. Budaghyan, Calderini, and Villa [1] introduced sufficient conditions for determining when CCZ-equivalence is strictly more general than EA-equivalence combined with taking inverses of permutations. On the algorithmic side, Kaleyski [9] surveyed invariants for EA- and CCZ-equivalence of APN and AB functions and evaluated their effectiveness in distinguishing inequivalent instances. Canteaut, Couvreur, and Perrin [2] presented an efficient algorithm for recovering the EA-equivalence transformation between quadratic functions, along with a new invariant based on the ortho-derivative. These works address the problem of testing whether two given functions are EA-equivalent.
@@
Both questions are intimately connected to the structure of EA-stabilizer groups. The EA-stabilizer of a function $F$ is the set of all pairs of affine permutations $(A_{\rm in},A_{\rm out})$ satisfying $A_{\rm out}\circ F\circ A_{\rm in}=F$; it measures how much “self-equivalence” the function possesses. By the orbit-stabilizer theorem, the size of each equivalence class is inversely proportional to the stabilizer size, which directly impacts both the total number of classes and the collision probability. However, explicit computation of stabilizers is infeasible except for very small dimensions, as the function space grows doubly exponentially with the input dimension.
@@ 1.1 Our Contributions
This paper resolves these questions by establishing that asymptotically almost all vectorial functions have trivial EA-stabilizers. Our main contributions are as follows:
@@
    Main theorem (Theorem 3.4): We prove that for a uniformly random vectorial function over $\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{m}$, the probability of having a nontrivial EA-stabilizer decays super-exponentially.
@@
    Counting equivalence classes (Corollary 3.7): We establish that the number of EA-equivalence classes is asymptotically equal to the total number of functions divided by the size of the EA-group, with vanishing relative error. This provides a simple and conceptually clear asymptotic formula grounded in the triviality of stabilizers.
@@
    Collision probabilities (Proposition 3.8 and Corollary 3.9): We derive upper bounds on collision probabilities for both EA- and CCZ-equivalence, and show that for EA-equivalence, the upper bound is asymptotically tight. This establishes that two independently sampled functions are EA-equivalent with super-exponentially small probability, validating random sampling strategies for cryptographic primitive design.
Collectively, these results demonstrate that functions with nontrivial EA-stabilizers form an exponentially rare subset. This complements existing approaches focused on specific cryptographic properties and provides a theoretical foundation for understanding the prevalence of nontrivial stabilizers in the space of vectorial functions.
@@ 1.2 Related Work
Previous work on stabilizers of Boolean functions has primarily focused on simpler group actions. Clausen [6] proved that almost all Boolean functions (not vectorial) have trivial stabilizers under the action of $\mathrm{GL}(n,2)$. Clote and Kranakis [7] established similar results for invariance groups under variable permutations. However, these results apply only to Boolean functions $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ and do not address EA-equivalence, which involves affine transformations on both domain and range, a key distinction for vectorial functions.
A different line of research focuses on counting equivalence classes. Hou [8] obtained exact formulas for affine equivalence classes and asymptotic formulas for EA-equivalence classes of Boolean functions using Burnside’s lemma and compound matrix rank estimates. Lu et al. [10] extended these results to vectorial Boolean functions, deriving exact formulas via Möbius inversion and asymptotic formulas by applying Hou’s matrix-theoretic techniques.
Our work unifies these two lines of research by establishing the connection between stabilizers and orbit enumeration for vectorial functions under EA-equivalence. We prove that almost all vectorial functions have trivial EA-stabilizers, extending the trivial-stabilizer results of Clausen and Clote-Kranakis to the vectorial EA setting. As an immediate consequence via the orbit-stabilizer theorem, we recover the asymptotic orbit counts of Hou [8] and Lu et al. [10] with an elementary probabilistic proof, and reveal the underlying structural reason: the naive estimate (total number of functions divided by the group size) is asymptotically exact because stabilizers are trivial with high probability.
@@ 1.3 Paper Organization
Section 2 reviews necessary background on group actions, boolean functions and EA-equivalence. Section 3 presents our main results, including the probabilistic analysis of EA-stabilizers and the asymptotic count of EA-classes. Section 4 concludes with a discussion of implications and open problems.
@@ 2.2 Vectorial Functions and Equivalence Relations
When $q=2$, functions in $\mathcal{F}$ are called vectorial Boolean functions. They play a central role in symmetric cryptography, where they are used as S-boxes in block ciphers and hash functions. The cryptographic quality of a vectorial function is measured by several properties. Differential uniformity quantifies resistance to differential cryptanalysis and is defined as the maximum number of solutions to F(x+a)F(x)=bF(x+a)-F(x)=b over all nonzero $a$ and all $b$. The Walsh spectrum and nonlinearity measure resistance to linear cryptanalysis, while the algebraic degree controls the complexity of algebraic attacks. Functions optimizing these properties, such as almost perfect nonlinear (APN) and almost bent (AB) functions, are of particular interest. Both EA-equivalence and CCZ-equivalence (defined below) preserve these cryptographic properties. For further details on these properties and their cryptanalytic significance, we refer to [5].
@@
$\Gamma:=\mathrm{AGL}(U)\times\mathrm{AGL}(W)$
acts on $\mathcal{F}$ by conjugating the input and output with affine maps:
$(A_{\rm out},A_{\rm in})\cdot F:=A_{\rm out}\circ F\circ A_{\rm in}.$
@@
$\bigl((P,a,Q,b)\cdot F\bigr)(x)=Q\,F(Px+a)+b.$
Two functions $F,G\in\mathcal{F}$ are EA-equivalent, denoted by $F\simeq_{\mathrm{EA}}G$, if they lie in the same $\Gamma$-orbit. The EA-stabilizer of $F$ is $\mathrm{Stab}_{\Gamma}(F)$.
@@
$\Gamma_{F}:=\{(x,F(x)):x\in U\}\subseteq U\times W=\mathbb{F}_{q}^{n+m}.$
@@
Two functions $F,G:U\to W$ are CCZ-equivalent (Carlet-Charpin-Zinoviev equivalent) if their graphs $\Gamma_{F}$ and $\Gamma_{G}$ are affine equivalent as subsets of $U\times W$. That is, $F$ and $G$ are CCZ-equivalent, denoted by $F\simeq_{\mathrm{CCZ}}G$, if there exists $L\in\mathrm{AGL}(U\times W)$ such that $L(\Gamma_{F})=\Gamma_{G}$.
CCZ-equivalence is more general than EA-equivalence: every pair of EA-equivalent functions is also CCZ-equivalent.
@@ 3.1 Fixed-Point and Orbit Structure of Affine Permutations
Let $\sigma$ be any permutation of a finite set of size $N$, with $f$ fixed points and $s$ orbits. Then
@@ 3.2 Probabilistic Analysis of EA-Stabilizers
Let $g=(P,a,Q,b)\in\Gamma$ be nontrivial. Then the number of $F\in\mathcal{F}$ satisfying $g\cdot F=F$ is bounded by
@@
$U=\bigsqcup_{i=1}^{s}O_{i},\qquad O_{i}=\{x_{i},\sigma(x_{i}),\sigma^{2}(x_{i}),\dots,\sigma^{L_{i}-1}(x_{i})\},$
where $L_{i}\geq 1$ is the length of orbit $O_{i}$.
Fix an orbit $O_{i}$ of length $L_{i}$. Iterating (1) starting from $x_{i}$ gives
$F(x_{i})$ $=Q\,F(\sigma(x_{i}))+b$
$=Q^{2}F(\sigma^{2}(x_{i}))+(I+Q)b$
@@
$=Q^{L_{i}}F(x_{i})+\sum_{j=0}^{L_{i}-1}Q^{j}b.$
Hence $F(x_{i})$ must satisfy
$\bigl(I-Q^{L_{i}}\bigr)F(x_{i})=\sum_{j=0}^{L_{i}-1}Q^{j}b.$ (2)
The number of solutions $F(x_{i})\in W$ to (2) is at most $q^{\dim\ker(I-Q^{L_{i}})}\leq q^{m}$. Once $F(x_{i})$ is chosen, the values $F(\sigma^{k}(x_{i}))$ for $1\leq k<L_{i}$ are uniquely determined by (1).
@@
Case 2: $(P,a)=(I,0)$ but $(Q,b)\neq(I,0)$. Now (1) becomes
$Q\,F(x)+b=F(x)\quad\Longleftrightarrow\quad(I-Q)F(x)=b\qquad(\forall x\in U).$
Since $(Q,b)\neq(I,0)$, either $Q\neq I$ or $b\neq 0$. If $Q\neq I$, then $\operatorname{rank}(I-Q)\geq 1$, so $\dim\ker(I-Q)\leq m-1$; the equation (IQ)F(x)=b(I-Q)F(x)=b has at most $q^{m-1}$ solutions for each $x$. If $Q=I$ and $b\neq 0$, there are no solutions. In either case, for each $x$ independently, $F(x)$ has at most $q^{m-1}$ choices. Hence
@@
Thus c=11/mc=1-1/m works when $m\geq 1$.
@@
Observe that $\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}$ if and only if there exists a nontrivial $g\in\Gamma$ such that $g\cdot F=F$. For each such $g$, define the event
$E_{g}:=\{F\in\mathcal{F}:g\cdot F=F\}=\mathrm{Fix}(g).$
By Lemma 3.3, we have
$\mathbb{P}(E_{g})=\frac{|\mathrm{Fix}(g)|}{|\mathcal{F}|}\leq q^{-(1-c)mq^{n}}.$
Applying the union bound over all $|\Gamma|-1$ nontrivial elements of $\Gamma$, we obtain
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)$ $=\mathbb{P}\Bigl(\bigcup_{g\in\Gamma\setminus\{\mathrm{id}\}}E_{g}\Bigr)$
$\leq\sum_{g\in\Gamma\setminus\{\mathrm{id}\}}\mathbb{P}(E_{g})$
$\leq(|\Gamma|-1)\cdot q^{-(1-c)mq^{n}}$
$\leq|\Gamma|\cdot q^{-(1-c)mq^{n}}.$
@@
so $|\mathrm{AGL}(U)|=q^{n}\cdot|\mathrm{GL}(n,q)|\leq q^{n^{2}+n}=q^{O(n^{2})}$. Similarly, $|\mathrm{AGL}(W)|=q^{O(m^{2})}$. Hence
$|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|\leq q^{O(n^{2}+m^{2})},$
@@
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{O(n^{2}+m^{2})}\cdot 2^{-(1/4)m2^{n}}\;=\;2^{-\Omega(m2^{n})}.$
@@
We evaluate the bound from Theorem 3.4 for parameters commonly used in symmetric cryptography ($q=2$, $c=3/4$).
  1. 1.
    44-bit S-boxes ($n=m=4$): The group size satisfies $\log_{2}|\Gamma|=\log_{2}|\mathrm{AGL}(4,2)|^{2}=2(4+\log_{2}|\mathrm{GL}(4,2)|)=2(4+\log_{2}20160)<2\cdot 19=38$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 4\cdot 16=16$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{38-16}=2^{22},$
    which does not yield a nontrivial bound. This is expected: the function space $|\mathcal{F}|=2^{64}$ is not large enough relative to the group size for asymptotic freeness to manifest.
  2. 2.
    88-bit S-boxes ($n=m=8$): We have $\log_{2}|\Gamma|\leq 2(8^{2}+8)=144$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 8\cdot 256=512$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{144-512}=2^{-368}.$
    This shows that a uniformly random $8$-bit S-box has a nontrivial EA-stabilizer with probability at most $2^{-368}$, an astronomically small quantity. The collision probability for EA-equivalence is similarly bounded by $2^{-mq^{n}+O(n^{2}+m^{2})}=2^{-2048+O(128)}\approx 2^{-1900}$.
Thus, for dimensions $n\geq 8$ (the regime relevant to modern block ciphers such as AES), the trivial-stabilizer property holds with overwhelming probability.
@@
The identity contributes $|\mathcal{F}|$. Each nontrivial $g$ contributes at most $q^{cmq^{n}}$ by Lemma 3.3. Hence
@@
$\frac{|\Gamma|\cdot q^{cmq^{n}}}{|\mathcal{F}|/|\Gamma|}=\frac{|\Gamma|^{2}\cdot q^{cmq^{n}}}{|\mathcal{F}|}=q^{-(1-c)mq^{n}+O(n^{2}+m^{2})}=o(1).$
To justify the exponent, note that $|\mathrm{AGL}(U)|=q^{n^{2}+O(n)}$ and $|\mathrm{AGL}(W)|=q^{m^{2}+O(m)}$, so $|\Gamma|=q^{n^{2}+m^{2}+O(n+m)}$, hence $|\Gamma|^{2}=q^{2n^{2}+2m^{2}+O(n+m)}=q^{O(n^{2}+m^{2})}$. Thus the relative error tends to zero. ∎
@@
$\frac{|\mathcal{F}/\Gamma|\cdot|\mathrm{AGL}(n,2)|\cdot|\mathrm{AGL}(m,2)|}{2^{m2^{n}}}=1+o(1).$
Consequently, the contribution of functions with nontrivial EA-stabilizers to the total count is exponentially negligible: their number is exponentially smaller than the full space. This observation is particularly relevant for structured searches in the design of cryptographic S-boxes.
@@
    $\mathbb{P}(F\sim_{\mathrm{EA}}G)\;\leq\;\frac{|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|}{|\mathcal{F}|}\;=\;q^{-mq^{n}+O(n^{2}+m^{2})}\;=\;q^{-\Omega(q^{n})}.$
@@
where $O_{1},\dots,O_{r}$ are the orbits. By the orbit-stabilizer theorem, the maximum orbit size is bounded by the group order. For EA-equivalence, $\max_{j}|O_{j}|\leq|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|$, and for CCZ-equivalence, $\max_{j}|O_{j}|\leq|\mathrm{AGL}(U\times W)|$. The size estimates follow from
$|\mathrm{AGL}(n,q)|=q^{n}\cdot|\mathrm{GL}(n,q)|=q^{n^{2}+O(n)}$
and $|\mathcal{F}|=q^{mq^{n}}$, giving the stated bounds. ∎
@@
$\mathbb{P}(F\sim_{\mathrm{EA}}G)=\frac{|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|}{|\mathcal{F}|}\,(1+o(1)).$

### 22 May 2026 v3

@@ 1 Introduction
Vectorial Boolean functions play a crucial role in symmetric cryptography, serving as S-boxes in block ciphers and hash functions. The classification of these functions is typically studied modulo certain equivalence relations that preserve important cryptographic properties. Extended-affine (EA) equivalence is one of the most fundamental such relations: it preserves differential uniformity, the Walsh spectrum, the nonlinearity, and the algebraic degree. A closely related but strictly coarser notion is CCZ-equivalence, introduced by Carlet, Charpin, and Zinoviev [4], which preserves differential uniformity, the Walsh spectrum, and the nonlinearity but does not in general preserve the algebraic degree [1].
@@
Both questions are intimately connected to the structure of EA-stabilizer groups. The EA-stabilizer of $F$ is the set of all triples $(A_{\rm in},A_{\rm out},L)$ satisfying $A_{\rm out}\circ F\circ A_{\rm in}+L=F$; it measures how much “self-equivalence” the function possesses. By the orbit-stabilizer theorem, the size of each equivalence class is inversely proportional to the stabilizer size, which directly impacts both the total number of classes and the collision probability. However, explicit computation of stabilizers is infeasible except for very small dimensions, as the function space grows doubly exponentially with the input dimension.
@@ 1.1 Our Contributions
The central contribution of this paper is to establish, for vectorial functions over finite fields, that the EA-group acts asymptotically freely: almost every function has trivial EA-stabilizer. From this single structural statement, both questions above are answered in tight asymptotic form, and a previously unknown matching lower bound on the EA-collision probability follows. The contributions are as follows.
@@
    Trivial EA-stabilizers (Theorem 3.4): for a uniformly random $F\in\mathcal{F}$, $\mathbb{P}(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\})=q^{-\Omega(mq^{n})}$. The proof is elementary: it bounds $|\mathrm{Fix}(g)|$ for each nontrivial $g\in\Gamma$ by analysing the orbit structure of affine permutations and applies a union bound. Restricted to ordinary (non-vectorial) Boolean functions and the action of $\mathrm{GL}(n,2)$, this recovers the trivial-symmetry theorems of Clausen [6] and Clote and Kranakis [7].
@@
    Probabilistic proof of the asymptotic EA-orbit count (Corollary 3.7): the number of EA-equivalence classes satisfies $|\mathcal{F}/\Gamma|=(|\mathcal{F}|/|\Gamma|)\,(1+o(1))$. The same asymptotic was proved analytically by Hou [8] (for $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$) and by Lu et al. [10] (for vectorial Boolean functions) using Burnside’s lemma combined with compound-matrix rank estimates and Möbius inversion. Our argument follows a fundamentally different route: the asymptotic is a one-line consequence of Theorem 3.4 via the orbit-stabilizer theorem, exposing the structural reason behind the naive estimate (trivial stabilizers with high probability) in place of analytic group-character computations.
@@
    Tight asymptotic collision probability (Proposition 3.8 and Corollary 3.9): we give upper bounds on the collision probability for both EA- and CCZ-equivalence which follow from orbit-stabilizer alone. Furthermore, by using Theorem 3.4, we prove a matching lower bound for EA-equivalence, yielding $\mathbb{P}(F\sim_{\mathrm{EA}}G)=(|\Gamma|/|\mathcal{F}|)(1+o(1))$. This sharp two-sided asymptotic for the EA-collision probability does not follow from Hou [8], Lu et al. [10], or earlier work, and is genuinely new.
@@ 1.2 Related Work
We organize prior work into three categories:
(i) Trivial-symmetry theorems for Boolean functions. Clausen [6] proved that almost all Boolean functions $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ have trivial $\mathrm{GL}(n,2)$-stabilizer; Clote and Kranakis [7] obtained an analogous statement for invariance under coordinate permutations. Both results are restricted to the codomain $\mathbb{F}_{2}$ and to a substantially smaller group than $\Gamma$. Theorem 3.4 extends this line to arbitrary vectorial codomains $\mathbb{F}_{q}^{m}$ and to the full EA-group, including the linear summand $L\in\mathrm{Hom}(U,W)$.
(ii) Counting EA-classes. Hou [8] obtained an asymptotic formula for the number of EA-equivalence classes of Boolean functions using Burnside’s lemma and compound-matrix rank estimates. Lu et al. [10] extended this to vectorial Boolean functions via Möbius inversion combined with Hou’s matrix-theoretic apparatus. Our Corollary 3.7 matches these asymptotics for vectorial Boolean functions and extends them to arbitrary $q$, with a shorter proof that does not require analytic estimates on compound matrices.
(iii) Structure of EA- and CCZ-equivalence. A complementary line of work studies the relationship between EA- and CCZ-equivalence and develops algorithms or invariants for distinguishing classes: Canteaut and Perrin [3] showed that CCZ-equivalence decomposes as EA-equivalence followed by an operation called twisting; Budaghyan, Calderini, and Villa [1] gave sufficient conditions for CCZ-equivalence to be strictly more general than EA-equivalence combined with inverses; Kaleyski [9] surveyed EA/CCZ invariants for APN and AB functions; and Canteaut, Couvreur, and Perrin [2] gave an efficient algorithm for recovering an EA-transformation between two quadratic functions. These works address testing EA-equivalence between explicit pairs of functions and are largely orthogonal to the present asymptotic-enumeration question.
@@ 1.3 Paper Organization
Section 2 reviews necessary background on group actions, Boolean functions, and EA-equivalence. Section 3 presents our main results, including the probabilistic analysis of EA-stabilizers and the asymptotic count of EA-classes. Section 4 concludes with a discussion of implications and open problems.
@@ 2.2 Vectorial Functions and Equivalence Relations
When $q=2$, functions in $\mathcal{F}$ are called vectorial Boolean functions. They play a central role in symmetric cryptography, where they are used as S-boxes in block ciphers and hash functions. The cryptographic quality of a vectorial function is measured by several properties. Differential uniformity quantifies resistance to differential cryptanalysis and is defined as the maximum number of solutions to F(x+a)F(x)=bF(x+a)-F(x)=b over all nonzero $a$ and all $b$. The Walsh spectrum and nonlinearity measure resistance to linear cryptanalysis, while the algebraic degree controls the complexity of algebraic attacks. Functions optimizing these properties, such as almost perfect nonlinear (APN) and almost bent (AB) functions, are of particular interest. EA-equivalence preserves all of the above properties, while CCZ-equivalence preserves differential uniformity, the Walsh spectrum, and nonlinearity, but does not in general preserve the algebraic degree [1]. For further details on these properties and their cryptanalytic significance, we refer to [5].
@@
$\Gamma:=\mathrm{AGL}(U)\times\mathrm{AGL}(W)\times\mathrm{Hom}(U,W)$
acts on $\mathcal{F}$ as follows: for $g=(A_{\rm in},A_{\rm out},L)\in\Gamma$ with $A_{\rm in}\in\mathrm{AGL}(U)$, $A_{\rm out}\in\mathrm{AGL}(W)$, and $L\in\mathrm{Hom}(U,W)$,
$(g\cdot F)(x):=A_{\rm out}\bigl(F(A_{\rm in}(x))\bigr)+L(x).$
@@
$\bigl((P,a,Q,b,L)\cdot F\bigr)(x)=Q\,F(Px+a)+Lx+b.$
Two functions $F,G\in\mathcal{F}$ are EA-equivalent, denoted by $F\simeq_{\mathrm{EA}}G$, if they lie in the same $\Gamma$-orbit. The EA-stabilizer of $F$ is $\mathrm{Stab}_{\Gamma}(F)$. The order of the EA group is
$|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|\cdot q^{nm}=q^{n^{2}+m^{2}+nm+O(n+m)}.$
The linear summand $L\in\mathrm{Hom}(U,W)$ in Definition 2.7 is essential: without it the action reduces to the (strictly smaller) action of $\mathrm{AGL}(U)\times\mathrm{AGL}(W)$ by pre- and post-composition, sometimes called affine equivalence [5].
@@
$\mathrm{Gr}(F):=\{(x,F(x)):x\in U\}\subseteq U\times W=\mathbb{F}_{q}^{n+m}.$
@@
Two functions $F,G:U\to W$ are CCZ-equivalent (Carlet-Charpin-Zinoviev equivalent) if their graphs $\mathrm{Gr}(F)$ and $\mathrm{Gr}(G)$ are affine equivalent as subsets of $U\times W$. That is, $F$ and $G$ are CCZ-equivalent, denoted by $F\simeq_{\mathrm{CCZ}}G$, if there exists $\Phi\in\mathrm{AGL}(U\times W)$ such that $\Phi(\mathrm{Gr}(F))=\mathrm{Gr}(G)$.
CCZ-equivalence is strictly more general than EA-equivalence: every pair of EA-equivalent functions is also CCZ-equivalent.
@@ 3.1 Fixed-Point and Orbit Structure of Affine Permutations
Let $\sigma$ be a permutation of a finite set $X$ of size $N$. A point $x\in X$ is a fixed point of $\sigma$ if $\sigma(x)=x$, and an orbit of $\sigma$ is an orbit of the cyclic group $\langle\sigma\rangle$ acting on $X$. Let $f$ denote the number of fixed points of $\sigma$ and $s$ the number of its orbits. Then
@@ 3.2 Probabilistic Analysis of EA-Stabilizers
Let $g=(P,a,Q,b,L)\in\Gamma$ be nontrivial, where $P\in\mathrm{GL}(n,q)$, $a\in U$, $Q\in\mathrm{GL}(m,q)$, $b\in W$, and $L\in\mathrm{Hom}(U,W)$. Then the number of $F\in\mathcal{F}$ satisfying $g\cdot F=F$ is bounded by
@@
$U=\bigsqcup_{i=1}^{s}O_{i},\qquad O_{i}=\{x_{i},\sigma(x_{i}),\sigma^{2}(x_{i}),\dots,\sigma^{\ell_{i}-1}(x_{i})\},$
where $\ell_{i}\geq 1$ is the length of orbit $O_{i}$.
Fix an orbit $O_{i}$ of length $\ell_{i}$. Iterating (1) starting from $x_{i}$ gives
$F(x_{i})$ $=Q\,F(\sigma(x_{i}))+Lx_{i}+b$
$=Q^{2}F(\sigma^{2}(x_{i}))+Q\bigl(L\sigma(x_{i})+b\bigr)+Lx_{i}+b$
@@
$=Q^{\ell_{i}}F(x_{i})+\sum_{j=0}^{\ell_{i}-1}Q^{j}\bigl(L\sigma^{j}(x_{i})+b\bigr).$
Setting $r_{i}:=\sum_{j=0}^{\ell_{i}-1}Q^{j}\bigl(L\sigma^{j}(x_{i})+b\bigr)\in W$ (which is determined once $g$ and the orbit representative $x_{i}$ are fixed), this implies
$\bigl(I-Q^{\ell_{i}}\bigr)F(x_{i})=r_{i}.$ (2)
The number of solutions $F(x_{i})\in W$ to (2) is at most $q^{\dim\ker(I-Q^{\ell_{i}})}\leq q^{m}$. Once $F(x_{i})$ is chosen, the values $F(\sigma^{k}(x_{i}))$ for $1\leq k<\ell_{i}$ are uniquely determined by (1).
@@
Case 2: $(P,a)=(I,0)$ but $(Q,b,L)\neq(I,0,0)$. Now (1) becomes
$Q\,F(x)+Lx+b=F(x)\quad\Longleftrightarrow\quad(I-Q)F(x)=Lx+b\qquad(\forall x\in U).$
If $Q\neq I$, then $\operatorname{rank}(I-Q)\geq 1$, so $\dim\ker(I-Q)\leq m-1$, and for each $x$ the equation (IQ)F(x)=Lx+b(I-Q)F(x)=Lx+b has at most $q^{m-1}$ solutions. Hence
@@
If $Q=I$, then the equation reduces to $Lx+b=0$ for all $x\in U$, which forces $L=0$ and $b=0$, contradicting nontriviality. So no $F$ satisfies the equation, and $|\mathrm{Fix}(g)|=0$. In either subcase, c=11/mc=1-1/m works when $m\geq 1$.
@@
[…]
$\{F\in\mathcal{F}:\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\}=\bigcup_{g\in\Gamma\setminus\{\mathrm{id}\}}\mathrm{Fix}(g),$
and by Lemma 3.3 each summand has probability $\mathbb{P}(\mathrm{Fix}(g))=|\mathrm{Fix}(g)|/|\mathcal{F}|\leq q^{-(1-c)mq^{n}}$. Applying the union bound over the $|\Gamma|-1$ nontrivial elements of $\Gamma$,
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)$ $\leq\sum_{g\in\Gamma\setminus\{\mathrm{id}\}}\mathbb{P}(\mathrm{Fix}(g))$
$\leq(|\Gamma|-1)\cdot q^{-(1-c)mq^{n}}\leq|\Gamma|\cdot q^{-(1-c)mq^{n}}.$
@@
so $|\mathrm{AGL}(U)|=q^{n}\cdot|\mathrm{GL}(n,q)|\leq q^{n^{2}+n}=q^{O(n^{2})}$. Similarly, $|\mathrm{AGL}(W)|=q^{O(m^{2})}$, and $|\mathrm{Hom}(U,W)|=q^{nm}$. Hence
$|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|\cdot|\mathrm{Hom}(U,W)|\leq q^{O(n^{2}+m^{2}+nm)}=q^{O((n+m)^{2})},$
@@
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{O((n+m)^{2})}\cdot 2^{-(1/4)m2^{n}}\;=\;2^{-\Omega(m2^{n})}.$
@@
We evaluate the bound from Theorem 3.4 for parameters commonly used in symmetric cryptography, i.e., $8$-bit S-boxes($n=m=8$): We have $\log_{2}|\Gamma|\leq 2(8^{2}+8)+64=208$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 8\cdot 256=512$. Hence
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{208-512}=2^{-304}.$
This shows that a uniformly random $8$-bit S-box has a nontrivial EA-stabilizer with probability at most $2^{-304}$, an astronomically small quantity Thus, for dimensions $n\geq 8$ (the regime relevant to modern block ciphers such as AES), the trivial-stabilizer property holds with overwhelming probability.
@@
The identity contributes $|\mathcal{F}|$. Each nontrivial $g$ contributes at most $q^{cmq^{n}}$ by Lemma $3.$ Hence
@@
$\frac{|\Gamma|\cdot q^{cmq^{n}}}{|\mathcal{F}|/|\Gamma|}=\frac{|\Gamma|^{2}\cdot q^{cmq^{n}}}{|\mathcal{F}|}=q^{-(1-c)mq^{n}+O((n+m)^{2})}=o(1).$
To justify the exponent, recall from the proof of Theorem 3.4 that $|\Gamma|=q^{O((n+m)^{2})}$, hence $|\Gamma|^{2}=q^{O((n+m)^{2})}$. Thus the relative error tends to zero. ∎
@@
$\frac{|\mathcal{F}/\Gamma|\cdot|\Gamma|}{2^{m2^{n}}}=1+o(1),$
where $|\Gamma|=|\mathrm{AGL}(n,2)|\cdot|\mathrm{AGL}(m,2)|\cdot 2^{nm}$. Consequently, the contribution of functions with nontrivial EA-stabilizers to the total count is exponentially negligible: their number is exponentially smaller than the full space. This observation is particularly relevant for structured searches in the design of cryptographic S-boxes.
@@
    $\mathbb{P}(F\sim_{\mathrm{EA}}G)\;\leq\;\frac{|\Gamma|}{|\mathcal{F}|}\;=\;q^{-mq^{n}+O((n+m)^{2})}\;=\;q^{-\Omega(q^{n})}.$
@@
where $O_{1},\dots,O_{r}$ are the orbits. By the orbit-stabilizer theorem, the maximum orbit size is bounded by the group order. For EA-equivalence, $\max_{j}|O_{j}|\leq|\Gamma|=q^{O((n+m)^{2})}$, and for CCZ-equivalence, $\max_{j}|O_{j}|\leq|\mathrm{AGL}(U\times W)|=q^{(n+m)^{2}+O(n+m)}$. Combined with $|\mathcal{F}|=q^{mq^{n}}$, this gives the stated bounds. ∎
We emphasise that these upper bounds rely only on the orbit-stabilizer theorem and the elementary estimates on the group orders; they do not require the trivial-stabilizer result of Theorem 3.4.
@@
$\mathbb{P}(F\sim_{\mathrm{EA}}G)=\frac{|\Gamma|}{|\mathcal{F}|}\,(1+o(1)).$